OpenStudy (anonymous):

Can you explain to me how to do this equation? Describe the locus of points whose coordinates satisfy both statements: x2 = 49 x2 + y2 = 85

OpenStudy (anonymous):

\[x =\pm \sqrt{7}\]

OpenStudy (anonymous):

What about \[x2 + y2 =85?\]

OpenStudy (anonymous):

\[\pm7, \pm6\]

OpenStudy (dumbcow):

replace 49 with x^2 49+y^2 = 85 y^2 = 85-49 =36 y = +- 6 solution is (+-7,+-6)

OpenStudy (anonymous):

x^2=49 so x^2+y^2=85 49+y^2=85, y^2=85-49 y=+-6 x=+-7

OpenStudy (anonymous):

Great, so then if i plot that, how would i find my radius? sorry to continue to ask questions! By the way thank you everyone for all of your help!

OpenStudy (anonymous):

\[7^2+y^2=85\]\[y^2=85-49\]\[y^2= \sqrt{36}\]\[y=\pm6\]

OpenStudy (anonymous):

how do i find the radius? or do i not need to?

OpenStudy (dumbcow):

x^2 +y^2 = r^2 therefore r^2 = 85 r = sqrt(85)

OpenStudy (anonymous):

THANK YOU! That makes complete sense!

OpenStudy (dumbcow):

your welcome, its the same for all circles (x-h)^2 + (y-k)^2 = r^2, (h,k) is center

OpenStudy (anonymous):

The locus is not a curve of a line. It is a set of four points (7,6), (-7,6), (-7,-6) and (7,-6) and they are where the lines x=+-7 or y=+-6 meets the circle x^2+y^2=85.

OpenStudy (anonymous):

I meant not a curve or line

OpenStudy (anonymous):

....lines meet the circle..